TANGENTS AND NORMALS 
find the points on the curve y=x^3 where the slope of the tangent is equal to x coordinate of the point.

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

Let the point be (a, b)
Slope = a
y = x^3
y' = 3x^2
a = 3a^2
a(3a-1) = 0
a = 0, \frac{1}{3}
Put in curve
a = 0\rightarrow b = 0
a = \frac{1}{3}\rightarrow b = \frac{1}{9}